OpenStudy (anonymous):

Solve. f(x)=2x^2-1/x ; x=sqrt2/2

OpenStudy (anonymous):

\[2x^2-1/\sqrt{x} ; x=\sqrt{2/2}\]

OpenStudy (anonymous):

1 -sqrt2

OpenStudy (anonymous):

that makes x=1

OpenStudy (anonymous):

oops\[x=\sqrt{2}/2\]

OpenStudy (anonymous):

Oh okay, then I think I must be right.

OpenStudy (anonymous):

well the book says the answer is \[2\sqrt{2}\]

OpenStudy (anonymous):

but i dont get that

OpenStudy (diyadiya):

\[2( \frac{1}{\sqrt{2}})^2- \frac{1}{1/\sqrt{2}}=1-\sqrt{2}\]

OpenStudy (anonymous):

well would help if i gave you guys the right equation

OpenStudy (anonymous):

\[2x^2-1/x\]

OpenStudy (diyadiya):

Yes still the answer is 1- sqqrt2

OpenStudy (anonymous):

\[(2(\sqrt{2}/2)^2-1)/\sqrt{2}/2\]

OpenStudy (anonymous):

thats what i need solved

OpenStudy (anonymous):

\[(4(\sqrt{2}/2)-2)/\sqrt{2}\]

OpenStudy (anonymous):

oops squared

OpenStudy (anonymous):

\[4(2/4)-2/\sqrt{2}\]

OpenStudy (anonymous):

which comes out to 0 for me

OpenStudy (anonymous):

so i assume i did something wrong

OpenStudy (diyadiya):

oh ok \[\frac{(2x)^2-1}{x}\]\[\frac{(2 \times \frac{1}{\sqrt{2}})^2-1}{ \frac{1}{\sqrt{2}}}\]\[(4 \times 1/2 \ -1)\sqrt{2}\]\[=\sqrt{2}\]

OpenStudy (anonymous):

its only the x squared

OpenStudy (anonymous):

not 2x

OpenStudy (diyadiya):

\[\frac{2(x)^2-1}{x}\]is this the equation?

OpenStudy (anonymous):

yes

OpenStudy (diyadiya):

It can't be because it gives zero unless zero is the answer

OpenStudy (anonymous):

thats what i was thinking

OpenStudy (anonymous):

but the relative minimum given by the book is

OpenStudy (anonymous):

\[(\sqrt{2}/2, 2\sqrt{2})\]

OpenStudy (anonymous):

which means the f(sqrt2/2) somehow is not 0

OpenStudy (diyadiya):

Wait \[x=\frac{\sqrt{2}}{2} \ or \ \frac{2}{\sqrt{2}}\]

OpenStudy (anonymous):

was 1/sqrt2 but rationalized is sqrt2/2

OpenStudy (diyadiya):

i'm getting the same answer sorry! once again check your question

OpenStudy (anonymous):

ok ill go through the entire problem with you

OpenStudy (diyadiya):

ok

OpenStudy (diyadiya):

wait

OpenStudy (anonymous):

i have to find the critical numbers of f(x)=2x+1/x using the first derivative test

OpenStudy (diyadiya):

Ok then its different

OpenStudy (anonymous):

so first critical number is 0 because f(x) is undefinedat 0

OpenStudy (anonymous):

then i find the derivative

OpenStudy (diyadiya):

Wait

OpenStudy (diyadiya):

i dont know much about this ..relative minimum

OpenStudy (anonymous):

oh im sure i got this part right

OpenStudy (anonymous):

2x/=1/x=(2x^2+1)/x

OpenStudy (anonymous):

first derivative

OpenStudy (anonymous):

(x(4x)-(2x^2+1))/X^2

OpenStudy (anonymous):

(4x^2-2x*2-1)/x^2

OpenStudy (anonymous):

(2x^2-1)/x^2

OpenStudy (anonymous):

thats the first derivative

OpenStudy (anonymous):

so 2x-1=0

OpenStudy (anonymous):

2x^2-1=0 oops

OpenStudy (anonymous):

x^2=1/2

OpenStudy (anonymous):

x=+/-1/sqrt2

OpenStudy (anonymous):

so critical numbers = 0,-sqrt2/2, and sqrt2/2

OpenStudy (anonymous):

0 is a vertical asymtope so its not a min or max

OpenStudy (anonymous):

but by the first derivative test

OpenStudy (anonymous):

i can see the interval ( -infinity, -sqrt2/2) is increasing which tells me -sqrt2/2 is the max

OpenStudy (diyadiya):

i don't understand this ..sorry But derivative of 2x+1/x=(2x^2-1)/x^2 right?

OpenStudy (anonymous):

the interval (sqrt2/2, infinty) is decreasing which means minimum

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

but since crit numbers = 0 or undefined

OpenStudy (anonymous):

you set the numerator to 0 to find crit numbers

OpenStudy (anonymous):

and since x=0 would make the denominator 0 that would be undefined

OpenStudy (anonymous):

hince the 3 critical numbers

OpenStudy (anonymous):

so this far i am correct

OpenStudy (anonymous):

now i need the y values for my max and min

OpenStudy (anonymous):

which means i plug my crit numbers into f(x)

OpenStudy (anonymous):

so i have to solve for f(sqrt2/2) and f(-sqrt2/2)

OpenStudy (anonymous):

f(x)= 2x+1/x

OpenStudy (diyadiya):

Ok so\[2x+ \frac{1}{x}= 2 \frac{1}{\sqrt{2}}+\frac{1}{1/\sqrt{2}}=\sqrt{2}+\sqrt{2}=2\sqrt{2}\]

OpenStudy (anonymous):

which equals (2x^2-1)/x

OpenStudy (anonymous):

well common denominator would be x

OpenStudy (anonymous):

so 2x(x)/x +1/x

OpenStudy (anonymous):

(2x^2+1)/x

OpenStudy (anonymous):

now plug in \[x=\sqrt{2}/2; x=-\sqrt{2}/2\]

OpenStudy (anonymous):

and the answers in the book i verified via cramster and calcchat.com

OpenStudy (diyadiya):

But in your question f(x) was \[2x+ \frac{1}{x}\]thats how you found derivative to be \[\frac{2x^2-1}{x^2}\] i'm sorry if i'm wrong i have no idea about this i'm just trying

OpenStudy (anonymous):

no those two are equal

OpenStudy (anonymous):

that just combining via common denominator

OpenStudy (anonymous):

oh wait that is the derivative

OpenStudy (anonymous):

but we plug the crit numbers back into f(x) to find y

OpenStudy (anonymous):

not the first derivative

OpenStudy (diyadiya):

Yes when we plug into f(x)\[2x+ \frac{1}{x}= 2 \frac{1}{\sqrt{2}}+\frac{1}{1/\sqrt{2}}=\sqrt{2}+\sqrt{2}=2\sqrt{2}\]

OpenStudy (callisto):

I'm so confused... Is f(x) = (2x^2 -1)/x or f(x) = (2x^2 +1)/x or f(x) = 2x^2 + (1/x) or f(x) = 2x^2 - (1/x)?

OpenStudy (anonymous):

and f(x) = (2x^2+1)/x

OpenStudy (anonymous):

or 2x+1/x

OpenStudy (anonymous):

both of those are the same

OpenStudy (diyadiya):

ok so when we plug in 1/sqrt2 to f(x) we get 2sqrt2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2(1/sqrt2)+1/1/sqrt2

OpenStudy (anonymous):

2/2sqrt +sqrt2

OpenStudy (anonymous):

2sqrt2/2+sqrt2

OpenStudy (anonymous):

=2sqrt2

OpenStudy (anonymous):

thats right thanks

OpenStudy (diyadiya):

welcome :)

OpenStudy (anonymous):

and now the negative

OpenStudy (anonymous):

must equal -2sqrt2

OpenStudy (callisto):

It's so messy here ... I dunno if this is sth you want 'cause i don't even know the question

OpenStudy (callisto):

(Nah.. seems problem's solved)